If one pipe takes 9 days and the other pipe 15 how long will bout take to fill the pool up
Answer by oberobic(2304)

Set up problems like these as fractions of the work done by each resource per unit of time.
.
Pipe 1 takes 9 days, so it does 1/9 of the job per day.
.
Pipe 2 takes 15 days, so it does 1/15 of the job per day.
.
Working together, how long does it take to do the job?
.
(1/9 + 1/15)*x = 1 job
.
(15+9)/(9*15)*x = 1
.
24/135 * x = 1
.
24*x = 135
.
x = 135/24 = 45/8 = 5 5/8 = 5.625 days
.
To check, determine how much each can do in that amount of time.
.
1/9*45/8 = 45/72 = 5/8
1/15 * 45/8 = 3/8
5/8 + 3/8 = 1
Correct
.
Working together will take 5 5/8 days.
.
Done.